# An ordering for the strength of functional dependence

### v1

@@ Abstract
We introduce a new dependence order that satisfies eight natural axioms that we propose for a global dependence order. Its minimal and maximal elements characterize independence and perfect dependence. Moreover, it characterizes conditional independence, satisfies information monotonicity, and exhibits several invariance properties. Consequently, it is an ordering for the strength of functional dependence of a random variable $Y$ on a random vector $\mathbb{X}\,.$ As we show, various dependence measures, such as Chatterjee’s rank correlation, are increasing in this order. We characterize our ordering by the Schur order and by the concordance order, and we verify it in models such as the additive error model, the multivariate normal distribution, and various copula-based models.
Keywords Chatterjee’s rank correlatifon, concordance order conditional independence, convex order, copula, dependence measure, dimension reduction, information gain inequality, rearrangements, perfect dependence, Schur order
@@ 1 Introduction and main results
In recent years, a large body of statistical literature has focused on dependence measures taking values in $[0,1]$, where 0 characterizes independence, and $1$ perfect functional dependence. A prominent example is Chatterjee’s rank correlation [7, 15] defined in Equation (2) below. While this measure gained popularity due to its fast nearest-neighbor–based estimator, it is also theoretically appealing due to further remarkable properties: it satisfies information monotonicity and characterizes conditional independence, enabling important applications such as model-free variable selection; see [16] for a survey. An open question concerns the identification of a dependence order that underlies the above type of dependence measures and describes their behavior across various models. In this paper, we fill this gap by introducing a new dependence order that allows us to answer the following question in the affirmative.
Is there a dependence order that characterizes independence through its minimal elements and perfect functional dependence through its maximal elements, satisfies information monotonicity and characterizes conditional independence, underlies Chatterjee’s rank correlation and related dependence measures, has a simple and interpretable form, can be verified in a variety of models, and applies to a large class of distributions?
@@
In contrast, for measures of regression dependence—which have become highly popular over the last decade—there is relatively little literature on underlying dependence orderings. Apart from our preliminary work [3, 5], research on global dependence orderings is conducted in [20, 22, 30, 54, 58, 59]. However, none of the dependence orderings considered in these references is able to describe, even in the simplest models, the behavior of dependence measures that characterize both independence and perfect functional dependence. For example, consider the additive error model
@@
where $\sigma\geq 0$ is a deterministic parameter scaling the noise $\varepsilon$ that is assumed to be independent of $f(\mathbb{X})\,.$ Of course, a suitable dependence measure quantifying the strength of dependence of $Y$ on $\mathbb{X}$ should be decreasing in $\sigma.$ Intuitively, one would attribute such a property to Chatterjee’s $\xi$; however, to the best of our knowledge, such ordering results for $\xi$ are known only under specific model assumptions, such as convolution semigroups. In the latter case, the desired monotonicity of $\xi$ in $\sigma$ is a direct consequence of the data processing inequality in [17]; see Remark 1.9(b) below.
In this paper, we introduce a new global dependence order—the conditional convex order—that describes the strength of functional dependence of a random variable $Y$ on a random vector $\mathbb{X}$. This ordering reflects the fundamental properties of a large class of dependence measures including $\xi$, and we verify it in settings like the additive error model (1), the multivariate normal distribution, and various copula-based models. Before we propose eight natural axioms for a global dependence ordering underlying $\xi$, recall the population version of Chatterjee’s rank correlation defined by
@@
see [7, 15]. It relies on the copula-based version in [22] and on the sensitivity measures in [12, 26]. Dependence measures like $\xi$ capture the variability of conditional distributions in various ways and have attracted a lot of attention in the past few years; see e.g. [6, 10, 13, 14, 21, 28, 29, 43, 60, 62, 67]. They attain values in the interval $[0,1]$ and have the zero-independence and max-functionality property, i.e., the minimal value 0 characterizes independence of $\mathbb{X}$ and $Y$, while the maximal value $1$ characterizes perfect (functional) dependence of $Y$ on $\mathbb{X}\,.$ The latter means that there exists a measurable function $f$ (not necessarily increasing or decreasing) such that $Y=f(\mathbb{X})$ almost surely.
For understanding the behavior of dependence measures in models such as (1), a suitable dependence order is essential. It is straightforward to see that $\xi$ is not increasing in the concordance order. For example, let $X$ and $Y$ be standard normal with $Y=\pm X.$ Then $(Y,X)$ is comonotone/countermonotone and thus maximal/minimal in the concordance order, but $\xi(Y,X)=1$ in both cases since $Y$ is a deterministic function of X.X\,.
1.1 Axioms for orderings of global dependence
We propose the following axioms for a global dependence ordering $\prec$ that compares the strength of functional dependence of $Y$ on $\mathbb{X}$ on a class $\mathcal{R}$ of random vectors. Well known axioms for positive dependence orderings are studied in [36, 37]. We denote by $\stackrel{{\scriptstyle\mathrm{d}}}{{=}}$ equality in distribution and write $\equiv$ if $\prec$ and $\succ$ holds true.
@@ 1.1 Axioms for orderings of global dependence
    Axiom 2 on reflexivity and transitivity is also considered in [20, 22, 58]. In contrast to positive dependence orderings, a suitable global dependence ordering $\prec$ cannot be expected to be antisymmetric (in the sense that $(Y,\mathbb{X})\equiv(Y^{\prime},\mathbb{X}^{\prime})$ implies $(Y,\mathbb{X})$ and $(Y^{\prime},\mathbb{X}^{\prime})$ have equal laws) because it is desirable that $(Y,Y)\equiv(Y,-Y)$ due to 4.
@@
    • [58, Axioms O4 and O5], according to which any bivariate random vector $(Y,X)$ is comparable with an independent version $(Y^{\perp},X^{\perp})$ and with a perfectly dependent version [^1] For a bivariate random vector $(Y,X)$, an independent resp. perfectly dependent version is a random vector $(Y^{\perp},X^{\perp})$ resp. $(Y^{\top},X^{\top})$ with $Y^{\perp}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\top}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y$ and $X^{\perp}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\top}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X$ such that $Y^{\top}$ and $X^{\top}$ are independent resp. $Y^{\top}$ is perfectly dependent on $X^{\top}$. An independent version exists if the underlying probability space is rich enough, for example, non-atomic. A perfectly dependent version exists if $\overline{\mathsf{Ran}(F_{Y})}\subseteq\overline{\mathsf{Ran}(F_{X})}$, where $\mathsf{Ran}(F_{Z})$ denotes the range of the distribution function of $Z$, and $\overline{A}$ is the closure of a set $A\subseteq[0,1]$. $(Y^{\top},X^{\top})$ from the same Fréchet class;
@@
    Axioms 5 and 6 state that additional information provided by further predictor variables increases the strength of functional dependence, and remains unchanged precisely when the additional information is irrelevant given the existing information. For dependence measures, information monotonicity is studied in [13]. On the level of dependence ordering, both axioms are new to the best of our knowledge. They yield a data processing inequality for $\prec$ in the sense that $(Y,\mathbb{Z})\prec(Y,\mathbb{X})$ whenever $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$. This implies $(Y,{\bf h}(\mathbb{X}))\prec(Y,\mathbb{X})$ for all measurable functions $\bf h$}.
@@
    Axiom 7 is stronger than the invariance properties in [20, Axiom O3] and [59, Axiom O3] where the function $g$ is assumed to be (increasing) linear. In [22], distributions from the same Fréchet class are transformed to the copula setting and no further invariance properties are studied.
@@
    Axiom 8 on distributional invariance implies that $(Y,\mathbb{X})\prec(Y^{\prime},\mathbb{X}^{\prime})$ only depends on the underlying copulas [^2] A $d$-variate copula is the distribution function of a random vector $(U_{1},\ldots,U_{d})$ with $U_{i}$ uniform on $(0,1)$ for all $i$. By Sklar’s theorem, every $d$-variate distribution function $F$ can be decomposed into $F=C\circ(F_{1},\ldots,F_{d}),$ where $F_{i}$ is the $i$th marginal distribution function and $C$ is a copula that is unique on $\mathsf{Ran}(F_{1})\times\cdots\times\mathsf{Ran}(F_{d})$. $C_{Y,\mathbb{X}}$ and $C_{Y^{\prime},\mathbb{X}^{\prime}}$ evaluated on the Cartesian products $\mathsf{Ran}(F_{Y})\times\mathsf{Ran}(F_{X_{1}})\times\cdots\mathsf{Ran}(F_{X_{p}})$ and $\mathsf{Ran}(F_{Y^{\prime}})\times\mathsf{Ran}(F_{X_{1}^{\prime}})\times\cdots\mathsf{Ran}(F_{X_{p^{\prime}}^{\prime}})$ of the respective marginal distribution functions. More specifically, since copulas are continuous, $(Y,\mathbb{X})\prec(Y^{\prime},\mathbb{X}^{\prime})$ only depends on the closed Cartesian products.
@@ 1.2 A new global dependence order
$\xi(Y,\mathbb{X})=\alpha\,\int_{\mathbb{R}^{p+1}}P(Y\geq y\mid\mathbb{X}=\mathbb{x})^{2}\;\mathrm{d}(P^{\mathbb{X}}\otimes P^{Y})(\mathbb{x},y)-\beta,$ (3)
@@
The first approach considers the variability of the conditional survival function $P(Y\geq y\mid\mathbb{X}=\mathbb{x})$ in $y$ for each fixed $\mathbb{x}\,.$ This might be seen as the more natural and more intuitive approach supported by a variety of literature on stochastic orderings for conditional distributions; see e.g. [11, 39, 51, 65, 66, 69]. In the setting of bivariate copulas, this approach is studied in [22, Theorem 2] where the dilation order and the dispersive order are used to construct regression dependence orders, which imply an ordering of $\xi$. However, as we discuss in Section 5.2, these orders are limited because they cannot be verified even in the simplest settings.
The second approach, which we explore in this paper, examines the variability of $P(Y\geq y\mid\mathbb{X}=\mathbb{x})$ in $\mathbb{x}$ for each fixed y.y\,. This perspective relates to the theory of random measures [34], and the works closest to our paper are [58, 59] where the variability of regression functions is studied. As it turns out, our approach is much more powerful as it yields an ordering that satisfies all the axioms above and thus reflects the fundamental properties of $\xi$. Further, as we show in Theorem 1.8 and Section 3, this order can be verified for many standard models, including the additive error model (1), the multivariate normal distribution, and various copula-based models.
Pursuing the second approach, we introduce below a new dependence order that we denote as conditional convex order. It compares conditional survival probabilities with respect to the convex order—the most common variability order studied in the literature; see [57, Section 3]. To this end, recall that, for bounded random variables $S$ and $T$, the convex order $S\leq_{cx}T$ is defined by $\mathbb{E}\varphi(S)\leq\mathbb{E}\varphi(T)$ for all convex functions $\varphi\colon\mathbb{R}\to\mathbb{R}$. We denote by $q_{Y}(v):=\inf\{y\mid F_{Y}(y)\geq v\},$ $v\in(0,1)$, the left-continuous quantile function of $Y$, also known as generalized inverse of $F_{Y}$.
@@
Before we present our main results, let us discuss why the conditional convex order is a suitable candidate for a global dependence order underlying $\xi$. First note that a necessary condition for $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ is $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$. This follows from the definition of the conditional convex order using that $S\leq_{cx}T$ implies $\mathbb{E}S=\mathbb{E}T$. We refer to this condition as marginal constraint. In this case, the denominator of $\xi$ in (2) is equal for $(Y,\mathbb{X})$ and $(Y^{\prime},\mathbb{X}^{\prime})$, and the conditional convex order only refers to the respective numerators.
Second, Chatterjee’s rank correlation is increasing in the conditional convex order: Since $S\leq_{cx}T$ implies $\mathrm{Var}(S)\leq\mathrm{Var}(T)$, it can easily be verified that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ yields $\xi(Y,\mathbb{X})\leq\xi(Y^{\prime},\mathbb{X}^{\prime})$. We provide this statement in a more general framework in Theorem 4.1.
Third, as studied in [13], convexity is essential for information monotonicity and, in particular, for the zero-independence and max-functionality property of dependence measures. It is straightforward to verify that minimal and maximal elements in conditional convex order characterize zero-independence and max-functionality: $Y$ and $\mathbb{X}$ are independent if and only if, for all $v\in(0,1)$, $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ is a.s. constant and thus minimal in convex order. Similarly, $Y$ perfectly depends on $\mathbb{X}$ if and only if, for all $v\in(0,1)$, $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ attains a.s. values in $\{0,1\}$ and thus is maximal in convex order within the class of $[0,1]$-valued random variables with mean $P(Y\geq q_{Y}(v))$. More generally, the conditional convex order allows a simple interpretation: little/large variability of the conditional probabilities $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ in convex order indicates low/strong dependence of $Y$ on $\mathbb{X}$—with the extreme cases of independence and perfect dependence discussed before.
@@
    To derive ordering results for models such as (1), we allow in Definition 1.2 a comparison of random vectors with different marginal distributions. If $Y$ and $Y^{\prime}$ have the same distribution, $(Y,\mathbb{X}),\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ simplifies to $P(Y\geq y\mid\mathbb{X})\leq_{cx}P(Y^{\prime}\geq y\mid\mathbb{X}^{\prime})$ for all $y\in\mathbb{R}$. For bivariate stochastically increasing distributions from the same Fréchet class, the conditional convex order coincides with the concordance order, as we show in Proposition 5.1. Recall that a bivariate distribution (function) is stochastically increasing (SI) if there exists a bivariate random vector $(W_{1},W_{2})$ with distribution (function) such that $W_{1}$ is SI in $W_{2}$, i.e., $\mathbb{E}[f(W_{1})\mid W_{2}=t]$ is increasing in $t$ for all increasing functions $f$ such that the expectations exist.
@@
    see Proposition A.4. The exceptional null sets in (4) are required due to possibly different continuity properties of the transformations $F_{Y}\circ q_{Y}$ and $F_{Y^{\prime}}\circ q_{Y^{\prime}}$ at the jump points of $F_{Y}$ and $F_{Y^{\prime}}$, respectively; see Example A.5 and Proposition A.1(ii) and (iii) for details. In contrast, the transformed survival functions $t\mapsto P(Y\geq q_{Y}(t))$ and $t\mapsto P(Y^{\prime}\geq q_{Y^{\prime}}(t))$ are left-continuous and, hence, they coincide under the marginal constraint for all $t$. This motivated us to define the conditional convex order through survival probabilities.

### v2

@@ Abstract
We introduce a new dependence order, termed the conditional convex order, whose minimal and maximal elements characterize independence and perfect dependence. Moreover, it characterizes conditional independence, satisfies information monotonicity, and exhibits several invariance properties. Consequently, it is an ordering for the strength of functional dependence of a random variable $Y$ on a random vector $\mathbb{X}\,.$ As we show, various recently studied dependence measures—including Chatterjee’s rank correlation, Wasserstein correlations, and rearranged dependence measures—are increasing in this order and inherit their fundamental properties from it. We characterize the conditional convex order by the Schur order and by the concordance order, and we verify it in settings such as additive error models, the multivariate normal distribution, and various copula-based models. Our results offer a unified perspective on the behavior of dependence measures across statistical models.
Keywords Chatterjee’s rank correlation; concordance order; conditional convex order; conditional independence; copula; dependence measure; dimension reduction; information monotonicity; optimal transport; perfect dependence; rearrangements; Schur order; Wasserstein correlation
@@ 1 Introduction and main results
A fundamental problem in regression analysis is to predict a response $Y$ given a random vector $\mathbb{X}$ of input variables. To construct parsimonious models, it is important to identify the relevant predictors. Recent literature has focused on model-free variable selection methods based on novel dependence measures $\kappa(Y,\mathbb{X})$ with the following remarkable properties [8, 17, 29, 62, 68]: $\kappa(Y,\mathbb{X})$ takes values in $[0,1]$, where 0 characterizes independence of $Y$ and $\mathbb{X}$ (zero-independence), and $1$ characterizes perfect dependence of $Y$ on $\mathbb{X}$ (max-functionality), i.e., there exists a measurable function $f$ such that $Y=f(\mathbb{X})$ almost surely. Moreover, $\kappa$ satisfies information monotonicity, that is, $\kappa(Y,\mathbb{X})\leq\kappa(Y,(\mathbb{X},\mathbb{Z}))$, and it characterizes conditional independence in the sense that $\kappa(Y,\mathbb{X})=\kappa(Y,(\mathbb{X},\mathbb{Z}))$ if and only if $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$. A prominent example is Chatterjee’s rank correlation $\xi$ defined in Equation (2) below. While the values 0 and $1$ have a clear interpretation, the intermediate values remain elusive. An open question concerns the identification of a dependence order that underlies the above type of dependence measures and ranks their values across various models. In this paper, we fill this gap by introducing a new dependence order that allows us to answer the following question in the affirmative.
Is there a dependence order that
  • characterizes independence through its minimal elements and perfect (functional) dependence through its maximal elements,
  • satisfies information monotonicity and characterizes conditional independence,
  • underlies Chatterjee’s rank correlation and related dependence measures,
  • has a simple and interpretable form,
  • can be verified in a variety of models, and
  • applies to a large class of distributions?
@@
In contrast, measures of regression dependence like $\xi$ behave fundamentally differently, since they range from independence to perfect (not necessarily monotone) dependence. Apart from our preliminary work [5, 4], research on global dependence orderings is conducted in [20, 22, 30, 56, 59, 60]. However, none of the dependence orderings in these references is able to describe, even in the simplest models, the behavior of dependence measures that characterize both independence and perfect dependence. For example, consider the additive error model
@@
where $\sigma\geq 0$ is a deterministic parameter scaling the noise $\varepsilon$ that is assumed to be independent of $f(\mathbb{X})\,.$ Since Chatterjee’s $\xi$ quantifies the strength of functional dependence [17], it should be decreasing in the noise parameter $\sigma$. Analogous monotonicity properties are expected to hold for related dependence measures such as Wasserstein correlations [68], rearranged dependence measures [62], kernel partial correlations [29], and measures of sensitivity [4]. However, to the best of our knowledge, general monotonicity properties for such dependence measures have not been studied in the literature so far.
In this paper, we resolve this issue and introduce a new global dependence order—the conditional convex order—that describes the strength of functional dependence of a random variable $Y$ on a random vector $\mathbb{X}$. This order reflects the fundamental properties of a large class of dependence measures, and we verify it in settings like the additive error model (1), the multivariate normal distribution, and various copula-based models. Before we propose eight natural axioms for a global dependence ordering underlying $\xi$, recall the population version of Chatterjee’s rank correlation defined by
@@
see [7, 16]. It relies on the copula-based version in [22] and on the sensitivity measures in [13, 26]. Dependence measures like $\xi$ capture the variability of conditional distributions in various ways and have attracted much attention in the past few years; see e.g. [6, 8, 12, 14, 15, 21, 28, 29, 42, 61, hörmann2025, 62, 68].
For understanding the behavior of dependence measures in various models, a suitable dependence order is essential. It is straightforward to see that $\xi$ is not increasing in the concordance order. For example, let $X$ and $Y$ be standard normal with $Y=\pm X.$ Then $(Y,X)$ is comonotone/countermonotone and thus maximal/minimal in the concordance order, but $\xi(Y,X)=1$ in both cases since $Y$ is a deterministic function of X.X\,.
1.1 Axioms for a global dependence ordering
We propose the following axioms for a global dependence ordering $\prec$ that compares the strength of functional dependence of $Y$ on $\mathbb{X}$ on a class $\mathcal{R}$ of random vectors defined on a common probability space $(\Omega,\mathcal{A},P)$. We denote by $\stackrel{{\scriptstyle\mathrm{d}}}{{=}}$ equality in distribution and write $\equiv$ if $\prec$ and $\succ$ holds true.
@@ 1.1 Axioms for a global dependence ordering
    Axiom 2 on reflexivity and transitivity is also considered in [20, 22, 59]. In contrast to positive dependence orderings, a suitable global dependence ordering $\prec$ cannot be expected to be antisymmetric (in the sense that $(Y,\mathbb{X})\equiv(Y^{\prime},\mathbb{X}^{\prime})$ implies $(Y,\mathbb{X})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(Y^{\prime},\mathbb{X}^{\prime})$) because it is desirable that $(Y,Y)\equiv(Y,-Y)$ due to Axiom 4.
@@
    • [59, Axioms O4 and O5], according to which any bivariate random vector $(Y,X)$ is comparable with an independent version $(Y^{\perp},X^{\perp})$ and with a perfectly dependent version [^1] For a bivariate random vector $(Y,X)$, an independent resp. perfectly dependent version is a random vector $(Y^{\perp},X^{\perp})$ resp. $(Y^{\top},X^{\top})$ with $Y^{\perp}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\top}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y$ and $X^{\perp}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\top}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X$ such that $Y^{\perp}$ and $X^{\perp}$ are independent resp. $Y^{\top}$ is perfectly dependent on $X^{\top}$. We generally assume that $(\Omega,\mathcal{A},P)$ is non-atomic and thus admits an independent and perfectly dependent version. $(Y^{\top},X^{\top})$ from the same Fréchet class;
@@
    Axioms 5 and 6 state that additional information provided by further predictor variables increases the strength of functional dependence, and remains unchanged precisely when the additional information is irrelevant given the existing information. For dependence measures, information monotonicity is studied in [12]. On the level of dependence orderings, both axioms are new to the best of our knowledge. They yield a data processing inequality for $\prec$ in the sense that $(Y,\mathbb{Z})\prec(Y,\mathbb{X})$ whenever $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$. This implies $(Y,{\bf h}(\mathbb{X}))\prec(Y,\mathbb{X})$ for all measurable $\bf h$}.
@@
    Axiom 7 is stronger than the invariance properties in [20, Axiom O3] and [60, Axiom O3] where $g$ is assumed to be (increasing) linear. In [22], distributions from the same Fréchet class are transformed to the copula setting and no further invariance properties are studied.
@@
    Axiom 8 on distributional invariance implies that $(Y,\mathbb{X})\prec(Y^{\prime},\mathbb{X}^{\prime})$ only depends on the underlying copulas [^2] A $d$-variate copula is the distribution function of a random vector $(U_{1},\ldots,U_{d})$ with $U_{i}\sim U(0,1)$ for all $i$. By Sklar’s theorem, every $d$-variate distribution function $F$ can be decomposed into $F=C\circ(F_{1},\ldots,F_{d}),$ where $F_{i}$ is the $i$th marginal distribution function and $C$ is a copula that is unique on $\mathsf{Ran}(F_{1})\times\cdots\times\mathsf{Ran}(F_{d})$. $C_{Y,\mathbb{X}}$ and $C_{Y^{\prime},\mathbb{X}^{\prime}}$ evaluated on the closed Cartesian products $\overline{\mathsf{Ran}(F_{Y})}\times\overline{\mathsf{Ran}(F_{X_{1}})}\times\cdots\overline{\mathsf{Ran}(F_{X_{p}})}$ and $\overline{\mathsf{Ran}(F_{Y^{\prime}})}\times\overline{\mathsf{Ran}(F_{X_{1}^{\prime}})}\times\cdots\overline{\mathsf{Ran}(F_{X_{p^{\prime}}^{\prime}})}$ of the respective marginal distribution functions.
@@ 1.2 A new global dependence order
$\xi(Y,\mathbb{X})=\alpha\,\int_{\mathbb{R}^{1+p}}P(Y\geq y\mid\mathbb{X}=\mathbb{x})^{2}\;\mathrm{d}(P^{Y}\otimes P^{\mathbb{X}})(y,\mathbb{x})-\beta,$ (3)
@@
The first approach considers the variability of the conditional survival probability $P(Y\geq y\mid\mathbb{X}=\mathbb{x})$ in $y$ for each fixed $\mathbb{x}\,.$ This might be seen as the more natural and more intuitive approach supported by a variety of literature on stochastic orderings for conditional distributions; [^3] For example, in optimal transport problems, optimal martingale couplings are based on a comparison of $P^{Y\mid X\in(-\infty,x]}$ in convex order for all $x\in\mathbb{R}$; see [10, Theorem 1.8] see [11, 38, 53, 66, 67]. In the setting of bivariate copulas, this approach is studied in [22] where the dilation order and the dispersive order are used to construct regression dependence orders, which imply an ordering of $\xi$. However, as we discuss in Section 5.2, these orders are limited because they cannot be verified even in the simplest settings.
The second approach, which we explore in this paper, examines the variability of $P(Y\geq y\mid\mathbb{X}=\mathbb{x})$ in $\mathbb{x}$ for each fixed y.y\,. This perspective relates to the theory of random measures [34], and the works closest to our paper are [59, 60] where the variability of regression functions is studied. As it turns out, our approach is much more powerful as it yields an ordering that satisfies all the axioms above and thus reflects the fundamental properties of $\xi$. Further, this order can be verified for various standard models as we show in Theorem 1.8 and Section 3.
Pursuing the second approach, we introduce below a new dependence order that we denote as conditional convex order. It compares conditional survival probabilities with respect to the convex order—the most common variability order studied in the literature; see [58, Section 3]. To this end, recall that, for bounded random variables $S$ and $T$, the convex order $S\leq_{cx}T$ is defined by $\mathbb{E}\varphi(S)\leq\mathbb{E}\varphi(T)$ for all convex functions $\varphi\colon\mathbb{R}\to\mathbb{R}$. We denote by $q_{Y}(v):=\inf\{y\mid F_{Y}(y)\geq v\},$ $v\in(0,1)$, the left-continuous quantile function of $Y$, also known as generalized inverse of $F_{Y}$.
@@
Before we present our main results, let us discuss why the conditional convex order is a suitable candidate for a global dependence order underlying $\xi$. First note that a necessary condition for $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ is $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$. This follows from the definition of the conditional convex order using that $S\leq_{cx}T$ implies $\mathbb{E}S=\mathbb{E}T$. We refer to this condition as marginal constraint. In this case, the denominator of $\xi$ in (2) is equal for $(Y,\mathbb{X})$ and $(Y^{\prime},\mathbb{X}^{\prime})$.
Second, the numerator of $\xi$ is increasing in the conditional convex order. This is a direct consequence of the fact that $S\leq_{cx}T$ implies $\mathrm{Var}(S)\leq\mathrm{Var}(T)$. It immediately follows that Chatterjee’s rank correlation is increasing in $\preccurlyeq_{ccx}$, i.e.,
$(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\quad\text{implies}\quad\xi(Y,\mathbb{X})\leq\xi(Y^{\prime},\mathbb{X}^{\prime}).$ (4)
We provide this statement in a more general framework in Theorem 4.1.
Third, as studied in [12], convexity is essential for information monotonicity and, in particular, for zero-independence and max-functionality of dependence measures. It is straightforward to verify that minimal and maximal elements in conditional convex order characterize zero-independence and max-functionality: On the one hand, $Y$ and $\mathbb{X}$ are independent if and only if, for all $v\in(0,1)$, $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ is a.s. constant and thus minimal in convex order. On the other hand, $Y$ perfectly depends on $\mathbb{X}$ if and only if, for all $v\in(0,1)$, $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ attains a.s. values in $\{0,1\}$ and thus is maximal in convex order within the class of $[0,1]$-valued random variables with mean $P(Y\geq q_{Y}(v))$. More generally, the conditional convex order allows a simple interpretation: Little/Large variability of the conditional probabilities $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ in convex order indicates low/strong dependence of $Y$ on $\mathbb{X}$—with the extreme cases of independence and perfect dependence discussed before.
@@
    To derive ordering results for models such as (1), we allow in Definition 1.2 a comparison of random vectors with different marginal distributions. If $Y$ and $Y^{\prime}$ have the same distribution, then $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ simplifies to $P(Y\geq y\mid\mathbb{X})\leq_{cx}P(Y^{\prime}\geq y\mid\mathbb{X}^{\prime})$ for all $y\in\mathbb{R}$. For bivariate stochastically increasing distributions from the same Fréchet class, the conditional convex order coincides with the concordance order, as we show in Proposition 5.1. Recall that a bivariate distribution (function) is stochastically increasing (SI) if there exists a bivariate random vector $(W_{1},W_{2})$ with this distribution (function) such that $W_{1}$ is SI in $W_{2}$, i.e., $\mathbb{E}[f(W_{1})\mid W_{2}=t]$ is increasing in $t$ for all increasing functions $f$ such that the expectations exist.
@@
    see Proposition A.4. The exceptional null set with respect to the Lebesgue measure $\lambda$ in (5) is required due to possibly different continuity properties of the transformations $F_{Y}\circ q_{Y}$ and $F_{Y^{\prime}}\circ q_{Y^{\prime}}$ at the jump points of $F_{Y}$ and $F_{Y^{\prime}}$, respectively; see Example A.5 and Proposition A.1iii and iv for details. In contrast, the transformed survival functions $t\mapsto P(Y\geq q_{Y}(t))$ and $t\mapsto P(Y^{\prime}\geq q_{Y^{\prime}}(t))$ are left-continuous and, hence, they coincide under the marginal constraint for all $t$. This motivated us to define the conditional convex order through survival probabilities.